find the height of a rectangulat prism with a surface area of 256 square meters. The sides of the base edges have a measure of 400 centimeters by 6 meters. give your solution in cm
Answer by addingup(3677)

400 * 600 = 240000 square cm
256 * 10000 = 2560000 square cm
---------------------------------------------------
2560000/240000 = 10.67 cm high